OpenStudy (anonymous):

how do I solve x-y=-6 6x-3y=-9 using systems of elemination?

OpenStudy (lukebluefive):

Rewriting the first equation as follows: \[y = x + 6\] Plugging this into the second equation gives: \[6x - 3(x + 6) = -9\] Multiplying: \[6x - 3x - 18 = -9\] Subtracting like terms: \[3x - 18 = -9\] Dividing both sides by 3: \[x - 6 = -3\] Adding 6 to both sides: \[x = 3\] Plugging back into our first derived equation: \[y = 3 + 6 = 9\]